OpenStudy (anonymous):

8th grade math...PLease helpp!! A company produces ink cartridges for printers. The setup cost for the run of blue cartridges is $7000 plus $8 per cartridges for materials. The company sells the cartridges for $24 each. How many cartridges must be produced and sold for the company to make a profit? PLEASE EXPLAIN OR SHOW WORK HOW YOU GOT UR ANSWER 8th grade math...PLease helpp!! A company produces ink cartridges for printers. The setup cost for the run of blue cartridges is $7000 plus $8 per cartridges for materials. The company sells the cartridges for $24 each. How many cartridges must be produced and sold for the company to make a profit? PLEASE EXPLAIN OR SHOW WORK HOW YOU GOT UR ANSWER @Mathematics

OpenStudy (anonymous):

437.5

OpenStudy (anonymous):

8x+7000=24X

OpenStudy (anonymous):

7000=16x

OpenStudy (anonymous):

Think of it like: PROFIT = REVENUE - COST How much money will they make? they will make 24 dollars for every cartridge sold. So if they number of cartridges sold was x, it would be:\[24x\] How mush will it cost to make those x cartridges? it will be 7000 start up fee (flat fee, only charged once), and then 8 dollars per cartridge. Thats:\[7000+8x\] Now lets set the equation equal to 0 to see how many cartridges they would have to make to break even:\[0=7000+8x-24x\iff -7000=-16x\iff \frac{7000}{16}=x=437.5\] But you cant make half a cartridge, so we have to make 438 to make profit.

OpenStudy (anonymous):

thankyou !

OpenStudy (lilg132):

438